If 0.65 mole of N2 occupies 18.0 L at a fixed particular temperature and pressure, how much volume would the gas occupy if the number of moles increased to 1.20?


PLEASE HELP


1
Expert's answer
2021-05-07T05:27:38-0400

pV = nRT

T = constant

p = constant

R = gas constant

"n_1 = 0.65 \;mol \\

V_1 = 18.0 \;L \\

pV_1 = n_1RT

\frac{V_1}{n_1} = \frac{RT}{p} \\

\frac{18.0}{0.65} = \frac{RT}{p} \\

27.69 = \frac{RT}{p} \\

\frac{V_2}{n_2} = \frac{RT}{p} \\

\frac{V_2}{n_2} = 27.69 \\

\frac{V_2}{1.20} = 27.69 \\

V_2 = 1.20 \times 27.69 = 33.23 \;L"

Answer: 33.23 L